OpenStudy (anonymous):

Earth has a diameter of approximately 7,920 miles and a volume that is roughly 49 times the volume of the moon. Find the diameter of the moon ?

OpenStudy (anonymous):

i cube root 1.01386 x 10^10 , right ?

OpenStudy (wolf1728):

Hold on I have to correct what I just posted

OpenStudy (wolf1728):

Sphere Volume = (4/3) * PI * r^3 Earth Volume = (4/3) * PI * 3,960^3 Earth Volume = 2.6012 x 10^11 cubic miles Moon volume = earth volume / 49 Moon volume = 2.6012 x 10^11 cubic miles / 49 Moon volume = 5.308576583 x 10^9 cubic miles 5.308576583 x 10^9 cubic miles = (4/3) * PI * r^3 r^3 = 5.308576583 x 10^9/[(4/3)* PI] r^3 = 1.267329305 x 10^9 r = 1082.1725 miles diameter = 1082.1725 miles times 2 = = 2,164.345 miles

OpenStudy (wolf1728):

Sorry I messed up the previous calculation. I'll delete it.

OpenStudy (anonymous):

;c its alright

OpenStudy (anonymous):

thx for trying to help

OpenStudy (wolf1728):

Actually, the calculations listed now are correct.